# Using Dynamic Safety Margins as Control Barrier Functions

### v3

@@ 5 DSM-CBF Synthesis Approach Summary
In this manuscript, we have shown that dynamic safety margins (DSMs), key components in the explicit reference governor ERG, are (vector-valued) CBFs for an augmented system obtained by concatenating the state and reference vectors. This insight allows the direct use of approaches that compute DSMs to synthesize CBFs. Existing ERG works like [23] outline different approaches to construct DSMs. This manuscript studied the Lyapunov-based approach, which requires finding a reference-dependent Lyapunov function (33) for the prestabilized system and evaluating the safety (34) and stability (36) threshold values. These ingredients can then be combined to form a DSM. Once a DSM is obtained, a safe constrained control policy that minimally adjusts the nominal controller can be readily implemented by solving (31). The following examples illustrate using the Lyapunov-based approach to construct DSM-CBFs and implementing (31). Additional examples can be found in [39], where the approach presented in this paper is specialized to Euler–Lagrange systems.
@@ 6.1 Anthill system
In the absence of inputs $u\!=\!0$, the system asymptotically tends to 00 if $|x_{0}|\!<\!1$ and diverges to infinity if $|x_{0}|>1$. Let $\mathcal{U}=[-u_{\max},u_{\max}]$ be the input constraint set, with $u_{\max}\!=\!\frac{4}{3\sqrt{3}}$.
@@
which satisfies the equilibrium conditions $\bar{x}(v)=v$ and $\bar{u}(v)=\pi\big{(}\bar{x}(v),v\big{)}=v-v^{3}$ for all $v\in\mathbb{R}$. The equilibrium point $\bar{x}(v)$ is attractive if $|x_{0}-v|<1$. This can be proven using the Lyapunov function
$V(x,v)=\int_{v}^{x}\!\!\!\!-f_{\pi}(x,v)dx=\frac{1}{2}(x-v)^{2}-\frac{1}{4}(x-v)^{4}.$ (43)
This anthill-shaped function (see Fig. 2) is such that, given the $\mathscr{K}_{\infty}$ functions $\gamma_{1}(c)=\frac{1}{4}c^{2}$ and $\gamma_{2}(c)=\frac{1}{2}c^{2}$, $\forall(x,v)\in\tilde{\mathcal{D}}=\{(x,v)\in\mathbb{R}\times\mathbb{R}:x\in(v-1,v+1)\}$,
@@
    DSM-CBF: We use $\eta=0.01$, the class $\mathscr{K}_{\infty}$ function $\alpha:c\mapsto 1.8c$, and $\rho:(v,r)\mapsto r-v$.
@@
    Candidate CBF: We use the functions $h_{1}(x)=x_{\max}-x$ and $h_{2}(x)=x_{\max}+x$ and tune their class $\mathscr{K}_{\infty}$ functions to illustrate the tradeoff between performance and feasibility that candidate CBFs suffer from.
@@
    and formulate the Backup CBF
    $h(x)=\min_{\tau\in[0,10]}x_{\max}-\Phi_{f_{\psi}}(x,\tau),$ (49)
@@
    where $\Phi_{f_{\psi}}(x_{0},\tau)$ is the solution of the backup dynamics $\dot{x}=f_{\psi}(x)\triangleq(x^{2}-1)x+\psi(x)$ at time $t=\tau$ with initial condition $x(0)=x_{0}$. The implementation is done following [21] with class $\mathscr{K}_{\infty}$ function $\alpha:c\mapsto 10c$.
@@
The simulation results for each approach are shown in Fig. 3. Both the ERG and the proposed DSM-CBF successfully enforce safety and reach the closest safe point to the desired reference but, the DSM-CBF exhibits performance that more closely resembles the nominal behavior. With enough tuning, the candidate CBFs were able to also maintain safety for the fist 10 seconds, albeit with poor performance and no guarantee they will remain feasible afterwards. This is unsurprising since the candidate CBFs were constructed without any understanding of the inherent control authority constraint. The backup CBF also fails in this example. Although the backup policy $\psi$ is safe for the starting condition considered, it loses control authority as soon as $x>v_{\max}$, which causes it to become infeasible. Conversely, the DSM-based CBF provided a systematic approach that successfully identified the constraints induced by the input saturation.
@@
Figure 3: Closed-loop behavior for the anthill system using each of the considered constrained control approaches. The dashed green and blue lines represent the evolution of the virtual reference $v(t)$ for the ERG and the DSM-CBF approaches, respectively. In addition, we show the performance of the candidate CBFs for different class $\mathscr{K}_{\infty}$ functions $\alpha_{1}=\alpha_{2}=\alpha:c\mapsto a_{i}c$, where $a_{i}\in\{7,0.25,0.15,0.1,0.07\}$. Larger $a_{i}$ corresponds to better input tracking, but also an earlier failure due to infeasibility, marked by a magenta ‘$\times$’ marker. The backup CBF is shown in orange and the ‘$\times$’ marker signals the point when the backup policy becomes unsafe, invalidating the approach.
@@ 6.2 Overhead Crane
$M(\boldsymbol{\mathbf{q}})\ddot{\boldsymbol{\mathbf{q}}}+V_{m}(\boldsymbol{\mathbf{q}},\dot{\boldsymbol{\mathbf{q}}})\dot{\boldsymbol{\mathbf{q}}}+G(\boldsymbol{\mathbf{q}})=Bu,$ (50)
@@
$M(\boldsymbol{\mathbf{q}})$ $=\begin{bmatrix}m_{c}+m_{p}&-m_{p}L\cos\theta\\ -m_{p}L\cos\theta&m_{p}L^{2}\end{bmatrix},\quad B=\begin{bmatrix}1\\ 0\end{bmatrix},$
$V_{m}(\boldsymbol{\mathbf{q}},\dot{\boldsymbol{\mathbf{q}}})$ $=\begin{bmatrix}0&m_{p}L\dot{\theta}\sin\theta\\ 0&0\end{bmatrix},\quad G(\boldsymbol{\mathbf{q}})=\begin{bmatrix}0\\ m_{p}gL\sin\theta\end{bmatrix},$
@@
where $k_{p}>0$, $k_{d}>0$ are the proportional and derivative gains. The equilibrium mapping is $\bar{x}:v\mapsto[v~{};0~{};0~{};0]$ and the reference-dependent Lyapunov function is
$V(\boldsymbol{\mathbf{x}},v)=\frac{1}{2}\dot{\boldsymbol{\mathbf{q}}}^{\top}M(\boldsymbol{\mathbf{q}})\dot{\boldsymbol{\mathbf{q}}}+m_{p}gL(1-\cos\theta)+\frac{1}{2}k_{p}(x-v)^{2},$
which satisfies Definition 4.21 with $\mathcal{D}_{v}=\mathbb{R}\times(-\pi,\pi)\times\mathbb{R}^{2}$ and also satisfies Assumption 4. We now define four types of constraints and, for each one, we find the associated safety threshold value or a continuously differentiable bounding function. For details on the construction of each $\Gamma^{*}_{i}(v)$, the reader is referred to [30, 39].
@@
    Payload constraints: $x+L\sin\theta\leq p_{\text{max}}$, with $p_{\text{max}}\geq x_{\text{max}}$. For this constraint, we were unable to find an analytical expression for $\Gamma_{5}^{*}(v)$ associated to $V$. However, as detailed in [30, Proposition 4], we can define
@@
For all simulations, we use linear class $\mathscr{K}_{\infty}$ functions given by $\alpha_{i}:c\mapsto\alpha_{i}c$ where $\alpha_{i}>0$ represents both the function and the scalar gain despite the slight abuse of notation. We choose $\alpha_{1}=\alpha_{2}=10$, $\alpha_{3}=20$, $\alpha_{4}=80$, $\alpha_{5}=2.2$ and $\alpha_{6}=80$. The navigation field we use is $\rho(v,r)=r-v$.
We compare with the following approaches.
@@
    Candidate CBF: For each constraint, we design a candidate CBF.
    1. 1)
      Position constraints: $h_{1}(\boldsymbol{\mathbf{x}})=\gamma_{1}(x-x_{\text{min}})+\dot{x}$ for $x_{\text{min}}\leq x$ and $h_{2}(\boldsymbol{\mathbf{x}})=\gamma_{2}(x_{\text{max}}-x)-\dot{x}$ for $x\leq x_{\text{max}}$.
    2. 2)
      Input constraints are added as upper and lower bounds for the optimization variable $-u_{\text{max}}\leq u\leq u_{\text{max}}$.
    3. 3)
      Angle constraints: $h_{3}(\boldsymbol{\mathbf{x}})=\gamma_{3}(\theta_{\text{max}}+\theta)+\dot{\theta}$ for $-\theta_{\text{max}}\leq\theta$ and $h_{4}(\boldsymbol{\mathbf{x}})=\gamma_{4}(\theta_{\text{max}}-\theta)-\dot{\theta}$ for $\theta\leq\theta_{\text{max}}$.
    4. 4)
      Payload constraints: $h_{5}(\boldsymbol{\mathbf{x}})=\gamma_{5}(p_{\text{max}}-x-L\sin\theta)-\dot{x}-L\dot{\theta}\cos\theta$ for $x+L\sin\theta\leq p_{\text{max}}$.
    Where $\gamma_{i}$ are positive constants. For all simulations, we use $\gamma_{i}=6$ and $\tilde{\alpha}_{i}=8,\forall i\in\{1,\ldots,4\}$, and $\gamma_{5}=4$, $\tilde{\alpha}_{5}=3$, where $\tilde{\alpha}_{i}:c\mapsto\tilde{\alpha}_{i}c$ is the class $\mathcal{K}_{\infty}$ function associated with $h_{i}$. It must be noted that all $h_{i}$ described above are candidate CBFs, meaning that they may not satisfy Definition 2 or have the control-sharing property given in Definition 2.4.
  4. (d)
@@
    $\psi(\boldsymbol{\mathbf{x}})=-u_{\max}.$ (54)
    For each constraint, we design a backup CBF and implement it according to [21]. For instance, the backup CBF associated to the payload constraint is
    $h(\boldsymbol{\mathbf{x}})\!=\!\min_{\tau\in[0,10]}p_{\max}\!-\!\boldsymbol{\mathbf{e}}_{1}^{\top}\Phi_{f_{\psi}}(\boldsymbol{\mathbf{x}},\tau)\!-\!L\sin\left(\boldsymbol{\mathbf{e}}_{2}^{\top}\Phi_{f_{\psi}}(\boldsymbol{\mathbf{x}},\tau)\right),$
    where $\boldsymbol{\mathbf{e}}_{i}$ is the $i$-th column of $I_{4}$ and $\Phi_{f_{\psi}}(\boldsymbol{\mathbf{x}}_{0},\tau)$ is the solution of the backup dynamics $f_{\psi}$ at time $t=\tau$ with initial condition $\boldsymbol{\mathbf{x}}(0)=\boldsymbol{\mathbf{x}}_{0}$. All class $\mathscr{K}_{\infty}$ functions are the same $\alpha(c)=300c$.
@@
$m_{c}$ (kg) $1$ $k_{p}$ $1$ $x_{\text{min}}$ (m) $-1.1$ $p_{\text{max}}$ (m) $1.1$
$m_{p}$ (kg) $0.5$ $k_{d}$ 0.1 $x_{\text{max}}$ (m) $1.1$ $\boldsymbol{\mathbf{x}}(0)$ (-) 00
$L$ (m) $0.7$ $k_{p,\kappa}$ $10$ $u_{\text{max}}$ (N) $4$ $v(0)$ (m) $0.1$
$\eta$ (-) 0.02 $k_{d,\kappa}$ $4$ $\theta_{\text{max}}$ (deg) $10$ $r$ (m) $1$
General parameters for the example are provided in Table 1. For the first simulation, shown in Fig. 5, we remove the angle constraint. In this case, all constrained control approaches achieve the safety objective. In terms of performance, the backup and candidate CBF approaches track the nominal behavior better than the proposed DSM-CBF approach, whereas the ERG exhibits the slowest response. The second simulation, shown in Fig. 6, includes the angle constraint $|\theta|\leq\theta_{\text{max}}$. The ERG and our proposed DSM-CBF approaches successfully enforce the constraints, with the proposed method achieving better performance. However, the candidate CBF-based controller becomes infeasible despite our best efforts to tune $\gamma_{i}$ and $\tilde{\alpha}_{i}$. This is unsurprising since candidate CBFs do not guarantee persistent feasibility. The backup policy $\psi$ is unsafe under the introduced angle constraints. This highlights that finding a suitable backup policy becomes more challenging as the number of constraints increases.
Figure 5: Closed-loop behavior in the absence of angle constraints. Even though the nominal controller violates the constraints, the ERG, candidate CBF, backup CBF and proposed DSM-CBF approaches all successfully enforce the safety requirements. The dashed blue and green lines are the virtual reference $v$ of the DSM-CBF and ERG approaches, respectively. Note that $v(t)$ may decrease in our approach, which is not possible in the ERG. This feature enlarges the feasible region of the CBF-based QP and allows the safe input $\beta(x,v)$ to be closer to the nominal input $\pi(x,r)$.
Figure 6: Closed-loop behavior in the presence of angle constraints. The ‘$\times$’ marks the instant in which the candidate and backup CBF policies become infeasible. The failure in the candidate CBF is likely due to the unavoidable backswing experienced by the payload angle $\theta(t)$. The failure in the backup CBF is due to the backup policy being unsafe with respect to the angle constraints. Conversely, the ERG and proposed DSM-CBF approaches successfully enforce the safety requirements.

### v4

@@ 5 DSM-CBF Synthesis Approach Summary
In this manuscript, we have shown that dynamic safety margins (DSMs), key components in the explicit reference governor (ERG), are (vector-valued) CBFs for an augmented system obtained by concatenating the state and reference vectors. This insight allows the direct use of approaches that compute DSMs to synthesize CBFs. Existing ERG works like [24] outline different approaches to construct DSMs. This manuscript studied the Lyapunov-based approach, which requires finding a reference-dependent Lyapunov function (33) for the prestabilized system and evaluating the safety (34) and stability (36) threshold values. These ingredients can then be combined to form a DSM. Once a DSM is obtained, a safe constrained control policy that minimally adjusts the nominal controller can be readily implemented by solving (31). The following examples illustrate using the Lyapunov-based approach to construct DSM-CBFs and implementing (31). Additional examples can be found in [42], where the approach presented in this paper is specialized to Euler–Lagrange systems.
@@ 6.1 Anthill system
In the absence of inputs u=0u\!=\!0, the system asymptotically tends to 0 if $|x_{0}|\!<\!1$ and diverges to infinity if $|x_{0}|>1$. Let $\mathcal{U}=[-u_{\max},u_{\max}]$ be the input constraint set, with $u_{\max}\!=\!\frac{4}{3\sqrt{3}}$.
@@
which satisfies the equilibrium conditions $\bar{x}(v)=v$ and $\bar{u}(v)=\pi\big(\bar{x}(v),v\big)=v-v^{3}$ for all $v\in\mathbb{R}$. The equilibrium point $\bar{x}(v)$ is attractive if $|x_{0}-v|<1$. This can be proven using a Lyapunov function constructed using the variable gradient method [37]
$V(x,v)=\int_{v}^{x}\!\!\!\!-f_{\pi}(\xi,v)\mathrm{d}\xi=\frac{1}{2}(x-v)^{2}-\frac{1}{4}(x-v)^{4}.$ (43)
This anthill-shaped function (see Fig. 2) is such that, given the class $\mathscr{K}$ functions $\gamma_{1}(c)=\frac{1}{4}c^{2}$ and $\gamma_{2}(c)=\frac{1}{2}c^{2}$, $\forall(x,v)\in\tilde{\mathcal{D}}=\{(x,v)\in\mathbb{R}\times\mathbb{R}:x\in(v-1,v+1)\}$,
@@
    DSM-CBF: We use $\eta=0.01$, the class $\mathscr{K}$ function $\alpha:c\mapsto 1.8c$, and $\rho:(v,r)\mapsto r-v$.
@@
    Candidate CBF: We use the functions $h_{1}(x)=x_{\max}+x$ and $h_{2}(x)=x_{\max}-x$ and tune their class $\mathscr{K}$ functions to illustrate the tradeoff between performance and feasibility that candidate CBFs suffer from.
@@
    which is a continuously differentiable approximation of $\psi^{*}(x)=-u_{\max}\text{sign}(x)$. The backup policy tries to drive the system to the origin using the maximum allowed input effort $u_{\max}$. Note that the set $\mathcal{S}^{0}=[-1,1]$ is forward invariant under the backup dynamics $\dot{x}=f_{\psi}(x)\triangleq(x^{2}-1)x+\psi(x)$. Now, we can formulate the backup CBF
    $h^{\psi}(x)=\min\left[\min_{\tau\in[0,T]}h^{\mathcal{C}}\big(\Phi_{f_{\psi}}(x,\tau)\big),h^{\mathcal{S}}\big(\Phi_{f_{\psi}}(x,T)\big)\right],$
@@
    where $T$ is the simulation horizon, $\Phi_{f_{\psi}}(x_{0},\tau)$ is the solution of the backup dynamics $f_{\pi}$ at time $\tau$ with initial condition $x(0)=x_{0}$, $h^{\mathcal{C}}(x)=x_{\max}-|x|$, and $h^{\mathcal{S}}(x)=1-|x|$. The implementation is done following [22] with class $\mathscr{K}$ function $\alpha:c\mapsto 7c$.
@@
The simulation results for each approach are shown in Fig. 3. All approaches are able to enforce safety and reach the closest safe point to the desired reference but differ in performance and computational complexity. As expected, our DSM-CBF approach exhibits behavior that more closely resembles the nominal behavior than the ERG. With enough tuning, the candidate CBFs were able to also maintain safety for the fist 10 seconds, albeit with poor performance and no guarantee they will remain feasible afterwards. This is unsurprising since the candidate CBFs were constructed without any understanding of the inherent control authority constraint. The backup CBF approach with horizon $T=1$ exhibits almost equal performance to our DSM-CBF approach. However, it must be noted that its computational footprint is significantly higher than our DSM-CBF approach due to the need at each time step to predict the backup trajectory forward in time. This computational footprint can be reduced by considering shorter time horizons (e.g. $T=0.1$), but performance worsens as showin in Fig. 3. The computational overhead of the safety filter associated with each appraoch is showin in Table 1. Moreover, the backup CBF approach is less numerically robust, as evidenced by the response to a longer time horizon $T=10$. The reason the backup CBF fails for longer horizons is that it is numerically searching for the open boundary of the domain of attraction $x<v_{\max}$. Since a slight overshoot in the numerical integration leads to the divergence of the backup trajectory prediction, the backup CBF fails if it pushes too close to the boundary of the domain of attraction. Conversely, the DSM-based CBF avoids this issue by using the stability threshold $0.99\,\overline{\Gamma}(v)$, which yields a closed set with enough numerical leeway that small perturbations do not exit the domain of attraction of the prestabilizing controller. Our DSM-based CBF provided a systematic approach that successfully identified the constraints induced by the input saturation.
Table 1: Safety Filter Compute Times
Anthill Crane
Approach Ave Max Ave Max
Nominal 0 0 0 0
ERG 0 0 0 0
Candidate CBF 4.81 5.02 - -
DSM-CBF (Proposed) 4.76 5.57 4.07 4.74
Backup CBF $T=0.1$ 5.95 6.02 9.04 10.12
Backup CBF $T=1.0$ 6.60 7.01 15.41 18.30
Backup CBF $T=5.0$ 15.03 20.16 24.51 32.62
@@
Figure 3: Closed-loop behavior for the anthill system using each of the considered constrained control approaches. The dashed green and blue lines represent the evolution of the virtual reference $v(t)$ for the ERG and the DSM-CBF approaches, respectively. In addition, we show the performance of the candidate CBFs for three different class $\mathscr{K}$ functions $\alpha_{1}=\alpha_{2}=\alpha:c\mapsto a_{i}c$, where $a_{i}\in\{7,0.15,0.07\}$. Larger $a_{i}$ corresponds to better input tracking, but also an earlier failure due to infeasibility, marked by a magenta $\times$’ marker. Finally, we show the backup CBF performance for three different horizons $T\in\{0.1,1,5\}$. The short horizon results in the conservative behavior and the long horizon induces numerical issues that ultimately cause the approach to fail, marked with the $\times$’ symbol. The horizon $T=1$ results in safe performance almost equal to our DSM-CBF.
@@ 6.2 Overhead Crane
$M(\boldsymbol{\mathbf{q}})\ddot{\boldsymbol{\mathbf{q}}}+C(\boldsymbol{\mathbf{q}},\dot{\boldsymbol{\mathbf{q}}})\dot{\boldsymbol{\mathbf{q}}}+G(\boldsymbol{\mathbf{q}})=Bu,$ (49)
@@
$M(\boldsymbol{\mathbf{q}})$ $=\begin{bmatrix}m_{c}+m_{p}&m_{p}L\cos\theta\\ m_{p}L\cos\theta&m_{p}L^{2}\end{bmatrix},\quad B=\begin{bmatrix}1\\ 0\end{bmatrix}$,
$C(\boldsymbol{\mathbf{q}},\dot{\boldsymbol{\mathbf{q}}})$ $=\begin{bmatrix}0&-m_{p}L\dot{\theta}\sin\theta\\ 0&0\end{bmatrix},\quad G(\boldsymbol{\mathbf{q}})=\begin{bmatrix}0\\ m_{p}gL\sin\theta\end{bmatrix}$,
@@
where $k_{p}>0$, $k_{d}>0$ are the proportional and derivative gains, and the equilibrium mapping is $\bar{x}:v\mapsto[v~;0~;0~;0]$. Since the overhead crane is an Euler–Lagrange system and $\pi$ is a passivity-based controller, let us formulate a storage function describing the total energy of the system [45]
$V(\boldsymbol{\mathbf{x}},v)=\frac{1}{2}\dot{\boldsymbol{\mathbf{q}}}^{\top}M(\boldsymbol{\mathbf{q}})\dot{\boldsymbol{\mathbf{q}}}+m_{p}gL(1-\cos\theta)+\frac{1}{2}k_{p}(x-v)^{2}.$
The storage function $V$ satisfies Definition 4.23 with $\mathcal{D}_{v}=\mathbb{R}\times(-\pi,\pi)\times\mathbb{R}^{2}$ and also satisfies Assumption 4. We now define four types of constraints and, for each one, we find the associated safety threshold value or a continuously differentiable bounding function. For details on the construction of each $\Gamma^{*}_{i}(v)$, the reader is referred to [32, 42].
@@
    Payload constraints: $x+L\sin\theta\leq p_{\text{max}}$, with $p_{\text{max}}\geq x_{\text{max}}$. For this constraint, we were unable to find an analytical expression for $\Gamma_{5}^{*}(v)$ associated to $V$. However, as detailed in [32, Proposition 4], we can define
@@
For all simulations, we use linear class $\mathscr{K}$ functions given by $\alpha_{i}:c\mapsto\alpha_{i}c$ where $\alpha_{i}>0$ represents both the function and the scalar gain despite the slight abuse of notation. We choose $\alpha_{i}=100$ for all $i\in\{1,\ldots,6\}$. The navigation field we use is $\rho(v,r)=r-v$.
For comparison, we consider the following approaches.
@@
    $\psi(\boldsymbol{\mathbf{x}})=u_{\max}\tanh\left(\frac{-K_{\psi}\boldsymbol{\mathbf{x}}}{u_{\max}}\right),$ (53)
    where $K_{\psi}=[3.15~-16.35~5.48~-5.16]$ was obtained linearizing the system around the origin and solving the LQR problem. The $\tanh$ term is a smooth clipping function that ensures $|\psi(\boldsymbol{\mathbf{x}})|\leq u_{\max}$. An invariant set associated with this backup policy is $\mathcal{S}^{0}=\{\boldsymbol{\mathbf{x}}\in\mathbb{R}^{4}\mid\boldsymbol{\mathbf{x}}^{\top}P_{\psi}\boldsymbol{\mathbf{x}}\leq\Gamma_{\psi}\}$, where $P_{\psi}$ is the symmetric, positive definite matrix that solves the Riccati equation associated with the LQR problem, and $\Gamma_{\psi}=2.5$ is determined heuristically through trial and error in the simulation. We formulate the backup CBF
    $h^{\psi}(\boldsymbol{\mathbf{x}})=\min\left[\min_{\tau\in[0,T]}h^{\mathcal{C}}\big(\Phi_{f_{\psi}}(\boldsymbol{\mathbf{x}},\tau)\big),h^{\mathcal{S}}\big(\Phi_{f_{\psi}}(\boldsymbol{\mathbf{x}},T)\big)\right],$
    where $T$ is the simulation horizon, $\Phi_{f_{\psi}}(\boldsymbol{\mathbf{x}}_{0},\tau)$ is the solution of the backup dynamics at time $\tau$ with initial condition $\boldsymbol{\mathbf{x}}(0)=\boldsymbol{\mathbf{x}}_{0}$,
    $h^{\mathcal{C}}(\boldsymbol{\mathbf{x}})=\min\begin{bmatrix}x_{\max}-|x|\\ \theta_{\max}-|\theta|\\ p_{\max}-x-L\sin\theta\end{bmatrix}$, (54)
    and $h^{\mathcal{S}}(\boldsymbol{\mathbf{x}})=\Gamma_{\psi}-\boldsymbol{\mathbf{x}}^{\top}P_{\psi}\boldsymbol{\mathbf{x}}$. The class $\mathscr{K}$ functions used are $\alpha(c)=40c$ for the position and payload constraints, $\alpha(c)=100c$ for the angle constraints and $\alpha(c)=400c$ for the terminal constraint $h^{\mathcal{S}}$.
@@
$m_{c}$ (kg) 11 $k_{p}$ 22 $x_{\text{min}}$ (m) 1.2-1.2 $p_{\text{max}}$ (m) $1.2$
$m_{p}$ (kg) $0.5$ $k_{d}$ 0.1 $x_{\text{max}}$ (m) $1.2$ $\boldsymbol{\mathbf{x}}(0)$ (-) 0
$L$ (m) $0.7$ $k_{p,\kappa}$ $10$ $u_{\text{max}}$ (N) 44 $v(0)$ (m) $0.1$
$\eta$ (-) 0.1 $k_{d,\kappa}$ 44 $\theta_{\text{max}}$ (deg) 2020 $r$ (m) 11
General parameters for the example are provided in Table 2, the performance of each approach is shown in Figure 5, and the computational burden of each approach is shown in Table 1. While all approaches achieve the goal and respect the safety constraints, their performance differs significantly. The ERG exhibits the slowest response but has the least computational burden. On the other hand, the backup CBF approach tracks the nominal behavior closely at the expense of high computational cost. Our DSM-CBF approach sits in the middle with a strict performance improvement over the ERG at a modest computational burden.
Figure 5: Closed-loop behavior of the overhead crane system under each of the considered constrained control approaches. The dashed green and blue lines represent the evolution of the virtual reference $v(t)$ for the ERG and DSM-CBF approaches, respectively. While only the trace for horizon $T=5$ for the backup CBF approach is shown, the computational burden of other horizon choices is shown in Table 1.